# Extremum Seeking with High-Order Lie Bracket Approximations: Achieving Exponential Decay Rate

### v2

@@ II-C Main idea
As follows from Lemma 2, the Lie bracket approximation approach provides a constructive solution to the extremum seeking problem which ensures the exponential convergence of the trajectories of system (4) to the optimal point in the case of a quadratic-like cost function, i.e., for $m=2$ in Assumption A2). However, for cases $m>2$, the above algorithm ensures only a polynomial decay rate. This can also be observed by analyzing the first-order Lie bracket system associated with system (4):
$\bar{x}=-\nabla J(\bar{x}).$
Note that the above first-order LBS is a complete average asymptote for (4), meaning that higher-order LBS approximations will be redundant [15]. Now, if for example $J=\dfrac{1}{2}(x-x^{*})^{2}$, then the Lie bracket system is linear, $\dot{\bar{x}}=-(\bar{x}-x^{*})$, and thus $x^{*}$ is its exponentially stable equilibrium point. In case $J=\dfrac{1}{4}(x-x^{*})^{4}$, the Lie bracket system takes the form $\dot{\bar{x}}=-(\bar{x}-x^{*})^{3}$, and its solutions are well-known to exhibit the polynomial decay rate $O(t^{-1/2})$ as $t\to\infty$ [30, 31]. Assume now that we can associate the properties of system (1) with a system which gives access to the third-order derivative of $J$, namely, with the system $\dot{\tilde{x}}=-J^{(3)}(\tilde{x})=-6(\tilde{x}-x^{*})$. Then the latter system turns out to be linear again, which, under certain assumptions, may imply the practical exponential stability of the extremum seeking system. A natural way of accessing higher order derivatives of the cost function is to excite the Lie brackets of corresponding order [14, 15]. For example, it is easy to see that
$[[[J,1],1],1](x)=-J^{(3)}(x).$
The main idea of this paper is to construct an extremum seeking system in the form
@@
so that, under a special choice of control vector fields $g_{k}$, $k=1,\dots,n_{u}\in\mathbb{N}$, dither signals $u_{k}^{\varepsilon}\in L^{\infty}_{[0,\varepsilon]}$, and a parameter $\varepsilon>0$, its trajectories approximate the trajectories of a system with high order Lie brackets:
@@
where $N\in\mathbb{N}$, $S_{\ell}\subset\{1,\dots,n_{u}\}^{\ell}$ denotes the sets of of multi-indices of the Lie brackets required for solving the extremum seeking problem, $g_{I_{\ell}}$ are the corresponding Lie brackets, and $c_{I_{\ell}}$ are constant parameters. For example, for system (3) with $r=2$, $f_{0}=0$, $f_{i}=g_{i}$, $i\in\{1,\dots,n\}$, we mean $N=2$, $S_{1}=\emptyset$, $S_{2}=\{(i,j):1\leq i<j\leq n\}$, $I_{2}:=(i,j)\in S_{2}$ $c_{I_{2}}=\nu_{ij}$.
One of the main tools exploited in this paper is the Chen–Fliess series expansion: under certain regularity assumptions on the control vector fields of system (5), the solutions of system (5) with $x(0)=x^{0}$ can be represented in the following form [32, 19]:
@@
×$\times\int\limits_{0}^{t}\int\limits_{0}^{s_{1}}...\int\limits_{0}^{s_{\ell-1}}u_{k_{1}}^{\varepsilon}(s_{1})\times\dots$
×$\qquad\qquad\qquad\times u_{k_{\ell}}^{\varepsilon}(s_{\ell})ds_{\ell}...ds_{1}+R(t),$
@@
$R(t)=$ $\sum_{k_{1},\dots,k_{N+1}=1}^{n_{u}}\int\limits_{0}^{t}\int\limits_{0}^{s_{1}}...\int\limits_{0}^{s_{N}}L_{g_{k_{N+1}}}\dots L_{g_{k_{2}}}g_{k_{1}}(x(s_{N+1}))$
×$\times u_{k_{1}}^{\varepsilon}(s_{1})\times\dots\times u_{k_{N+1}}^{\varepsilon}(s_{N+1})ds_{N+1}...ds_{1}.$
@@ III-A Two-input extremum seeking system
To simplify the presentation, in this section, we assume $x\in\mathbb{R}$. To steer the solutions of an extremum seeking system towards the direction of high-order Lie brackets, we refer to control approaches from nonholonomic systems theory [9, 10, 33, 17, 34, 35, 19, 20]. We focus here on the two-input systems of form (5):
@@
Suppose also that $u_{1}^{\varepsilon}(t)=\varepsilon^{1/N-1}v_{1}(t/\varepsilon)$, $u_{2}^{\varepsilon}(t)={\varepsilon^{1/N-1}}v_{2}(t/\varepsilon)$, with the dithers $v_{1}(t/\varepsilon),v_{2}(t/\varepsilon)$ exciting the Lie bracket
$g_{I_{N}}(z)=\big{[}\big{[}\dots[g_{1},\underbrace{g_{2}],g_{2}\big{]},\dots,g_{2}}_{N-1\text{ times}}]\dots]\big{]}(z)$
at time $t=\varepsilon$, with $I_{N}=(1,\underbrace{2,\dots,2}_{N-1\text{ times}})$, in the sense that the Chen–Fliess series expansion (7) takes the form
@@
and all the other Lie brackets of length from 1 to $N$ do not appear in the above expansion. One way to construct such inputs is described in [34, 35], other approaches can be found in, e.g., [9, 10, 15]. The concrete examples are as follows, as mentioned in [20].
@@
  • the inputs $v_{1}(t/\varepsilon)=v_{1}^{12}(t/\varepsilon)=2\sqrt{\kappa_{12}\pi}\cos\left({2\kappa_{12}\pi t}/{\varepsilon}\right)$,
    $v_{2}(t/\varepsilon)=v_{2}^{12}(t/\varepsilon)=2\sqrt{{\kappa_{12}\pi}}\sin\left({2\kappa_{12}\pi t}/{\varepsilon}\right)$, $\kappa_{12}\in\mathbb{Z}$,
    excite the first order Lie bracket $[g_{1},g_{2}]$;
  • the inputs
    $v_{1}(t/\varepsilon)=v_{1}^{122}(t/\varepsilon)=-2\left({4\kappa_{122}\pi}\right)^{\frac{2}{3}}\cos\left({4\kappa_{122}\pi t}/{\varepsilon}\right)$,
    $v_{2}(t/\varepsilon)=v_{2}^{122}(t/\varepsilon)=\left({4\kappa_{122}\pi}\right)^{\frac{2}{3}}|\cos\left({2\kappa_{122}\pi t}/{\varepsilon}\right)$,
    $\kappa_{112}\in\mathbb{Z}$, excite the second order Lie bracket $[[g_{1},g_{2}],g_{2}]$;
  • the inputs
    $v_{1}(t/\varepsilon)=v_{1}^{1222}(t/\varepsilon)=6\left(2\kappa_{1222}\pi\right)^{\frac{3}{4}}\sin\left({6\kappa_{1222}\pi t}/{\varepsilon}\right)$,
    $v_{2}(t/\varepsilon)=v_{2}^{1222}(t/\varepsilon)=2\left(2\kappa_{1222}\pi\right)^{\frac{3}{4}}\cos\left({2\kappa_{1222}\pi t}/{\varepsilon}\right)$,
@@
Remark 1
By examining (3) in light of the design choices in Statement 1 above, we note:
  • the choice of $v_{1}^{12}$ and $v_{2}^{12}$ to excite $[g_{1},g_{2}]$ fits the condition provided in [15, Section 4.1] when $p_{1}+p_{2}=1/2+1/2=1$, which guarantees bounded first-order LBS as $\varepsilon\rightarrow 0$ and that the realized first-order LBS is a complete average asymptote (i.e., higher-order LBS approximations are made redundant by that design in that they vanish as $\varepsilon\rightarrow 0$).
  • the choices of $v_{1}^{122}$ and $v_{2}^{122}$ to excite $[[g_{1},g_{2}],g_{2}]$ fits the condition provided in [15, Section 4.1] when $p_{1}+p_{2}+p_{2}=2/3+2/3+2/3=2$, which means that the second-order LBS is designed as the complete average asymptote. Also, due to the design choice of the dither signals with $k_{1}=4\kappa_{122}$ and $k_{2}=2\kappa_{122}$, coefficients resulting from the iterated integrals for first-order terms and other second-order terms (except the excited bracket) vanish. This also leads to a bounded second-order LBS as $\varepsilon\rightarrow 0$.
  • the choices of $v_{1}^{1222}$ and $v_{2}^{1222}$ to excite $\big{[}[[g_{1},g_{2}],g_{2}],g_{2}\big{]}$ fits the condition provided in [15, Section 4.1] when $p_{1}+p_{2}+p_{2}+p_{2}=3/4+3/4+3/4+3/4=3$, which means that the third-order LBS is designed as the complete average asymptote. Also, due to the design choice of the dither signals with $k_{1}=6\kappa_{1222}$ and $k_{2}=2\kappa_{1222}$, coefficients resulting from the iterated integrals for first-order terms, second-order terms and other third-order terms (except the excited bracket) vanish. This also leads to a bounded third-order LBS as $\varepsilon\rightarrow 0$.
@@
$[g_{1}\circ J,g_{2}\circ J](x)=-\varphi(J(x))\nabla J(x),$ (11)
@@
The argumentation is similar to the proof of the practical exponential stability in [13, Theorem 3], so we only explain here how to derive exponential decay rate to a neighborhood of $x^{*}$ based on the representation (9). Given any $\delta>0$, let $D^{\prime}$ be any compact set such that $\overline{B_{\delta}(x^{*})}\subset D^{\prime}\subset D$. Similarly to the proof ofn [13, Theorem 3], one can show that there exists an $\varepsilon_{0}>0$ such that solutions of system (8) with $x^{0}\in B_{\delta}(x^{*})$ and $\varepsilon\in(0,\varepsilon_{0})$ are well-defined in $D^{\prime}$ for all $t\in[0,\varepsilon]$. To estimate the remainder in (12), denote
$c_{Fm}=\sup\limits_{x\in D^{\prime}}\sum_{k_{1},\dots,k_{m+1}=1}^{2}L_{g_{k_{m}}}\dots L_{g_{k_{2}}}g_{k_{1}}\circ J(x),$
$c_{u}=\max\limits_{k=1,2}\sup\limits_{t\in[0,\varepsilon]}|v_{k}(t/\varepsilon)|$. Then it is easy to see that
$\|R(\varepsilon)\|=$
$\sum_{k_{1},\dots,k_{m+1}=1}^{2}$ $\int\limits_{0}^{t}\int\limits_{0}^{s_{1}}...\int\limits_{0}^{s_{m}}L_{g_{k_{m+1}}}\dots L_{g_{k_{2}}}g_{k_{1}}\circ J(x(s_{m+1}))$
×$\times u_{k_{1}}^{\varepsilon}(s_{1})\times\dots\times u_{k_{m+1}}^{\varepsilon}(s_{m})ds_{m}...ds_{1}$
$\leq$ $c_{R}\varepsilon^{1+\frac{1}{m}},$
where $c_{R}=c_{u}^{m+1}c_{Fm}$. The above estimate together with the representation (12), Assumption 2, and triangular inequality, implies that
$\|x(\varepsilon)-x^{*}\|^{2}\leq$ $\|x^{0}-x^{*}\|^{2}(1-2\varepsilon\gamma_{m})+\varepsilon^{1+\frac{1}{m}}\sigma,$
with $\gamma_{m}=c_{m}(\beta_{1}-\varepsilon c_{m}\beta_{2}^{2})$, $\sigma=c_{R}^{2}\varepsilon^{1+\frac{1}{m}}+2c_{R}\delta(1+\varepsilon c_{m}\beta_{2})$. Then following the argumentation of [13, Steps 3.I-4.I], we can establish the practical exponential decay rate to an arbitrary small neighborhood of $x^{*}$. ∎
@@
As in the paper [13], it is also possible to relax regularity assumption on the control vector fields. Namely, requirement of $g_{k}\circ J$, $k=1,2$, being $m$ times continuously differentiable in $D$ can be replaced with the following: $g_{k}\circ J\in C^{m}(D\setminus\{x^{*}\});\mathbb{R}$ and $L_{g_{k_{N+1}}}\dots L_{g_{k_{2}}}g_{k_{1}}\circ J\in C(D;\mathbb{R})$ for all $N\in\{1,\dots,m\}$, $k_{1},\dots,k_{m+1}\in\{1,2\}$. This relaxation is particularly important for deriving conditions for the “classical” exponential stability in the sense of Lyapunov, meaning that the trajectories of the extremum seeking system converge to the point $x^{*}$, rather than merely to its neighborhood. Another important condition for achieving classical exponential stability is the property of vanishing amplitudes, which requires that $g_{k}\circ J\to 0$ as $x\to x^{*}$ (see [13, Theorem 3, Part II]). Since selecting such vector fields becomes increasingly challenging in the case of higher-order Lie brackets, we leave this task, along with a rigorous formulation of the corresponding exponential stability properties, for furture research.
@@ III-B Alternative design approaches
with $u_{k}^{\varepsilon}(t)=\varepsilon^{-3/4}v_{k}^{\varepsilon}(t)$, $k=1,2,3$, with $v_{k}(t/\varepsilon)$ exciting the Lie bracket $g_{(1,2,3)}=[[g_{1},g_{2}],g_{3}]$, and $g_{1},g_{2},g_{3}$ such that
@@
Let $\varphi_{2}\in C^{1}(\mathbb{R};\mathbb{R})$ be any given function, the functions $g_{1},g_{2}\in C^{1}(\mathbb{R};\mathbb{R})$ satisfy (11) with $\varphi=\varphi_{2}$, and let $g_{3}=-\varphi_{2}$. Then
$\big{[}[g_{1}\circ J,g_{2}\circ J],g_{3}\circ J\big{]}(x)=-\varphi_{2}^{2}(x)J^{(2)}(x)\text{ for all }x\in D.$
Proof:
Indeed, under the above conditions,
[…]
$=\big{[}\varphi_{2}(x)\nabla J(x),\varphi_{2}(J(x))\big{]}(x)=-\varphi_{2}^{2}(J(x))J^{(2)}(x).$
@@
With the exception of special cases (e.g., Newton-based ES as in [14]), in general, the even-order derivatives are rather not helpful for classic extremum seeking problems. The goal of considering this case here is to illustrate alternative approaches to designing extremum seeking systems, both in terms of selecting control vector fields and determining the number of dithers. The formula presented in Lemma 3 can be easily extended to higher-order scenarios. For example, given a four inputs system (5) with $n_{u}=4$ and any function $\varphi_{3}\in C^{1}(\mathbb{R};\mathbb{R}^{+})$ such that $\sqrt{\varphi_{3}}\in C^{1}(\mathbb{R}^{+};\mathbb{R}^{+})$, let the functions $g_{1},g_{2},g_{3}$ be chosen as in Lemma 3 with $\varphi_{2}=\sqrt{\varphi_{3}}$. Then $\big{[}\big{[}[g_{1}\circ J,g_{2}\circ J],g_{3}\circ J\big{]},g_{4}\circ J\big{]}(x)=-\varphi_{3}^{2}J^{(3)}(x)$.
@@
The proof mostly repeats the proof of Theorem 1.

### v3

@@ II-C Main idea
As follows from Lemma 2, the Lie bracket approximation approach provides a constructive solution to the extremum seeking problem which ensures the exponential convergence of the trajectories of system (3) to the optimal point in the case of a quadratic-like cost function, i.e., for $m=2$ in Assumption A2). However, for cases $m>2$, the above algorithm ensures only a polynomial decay rate. This can also be observed by analyzing the first-order Lie bracket system associated with system (3): $\dot{\bar{x}}=-\nabla J(\bar{x}).$ Note that the above first-order LBS is a complete average asymptote for (3), meaning that higher-order LBS approximations will be redundant [13]. Now, if for example $J=\tfrac{1}{2}(x-x^{*})^{2}$, then the Lie bracket system is linear, $\dot{\bar{x}}=-(\bar{x}-x^{*})$, and thus $x^{*}$ is its exponentially stable equilibrium point. In case $J=\tfrac{1}{4}(x-x^{*})^{4}$, the Lie bracket system takes the form $\dot{\bar{x}}=-(\bar{x}-x^{*})^{3}$, and its solutions are well-known to exhibit the polynomial decay rate $O(t^{-1/2})$ as $t\to\infty$ [26, 27]. Assume now that we can associate the properties of system (1) with a system which gives access to the third-order derivative of $J$, namely, with the system $\dot{\tilde{x}}=-J^{(3)}(\tilde{x})=-6(\tilde{x}-x^{*})$. Then the latter system turns out to be linear again, which, under certain assumptions, may imply the practical exponential stability of the extremum seeking system. A natural way of accessing higher order derivatives of the cost function is to excite the Lie brackets of corresponding order [12, 13]. For example, it is easy to see that $J,1],1],1](x)=-J^{(3)}(x).$ The main idea of this paper is to construct an extremum seeking system
@@
so that, under a special choice of vector fields $g_{k}$, dither signals $u_{k}^{\varepsilon}{\in}L^{\infty}_{[0,\varepsilon]}$, and an $\varepsilon{>}0$, its trajectories approximate the trajectories of a system with high order Lie brackets:
@@
where $N\in\mathbb{N}$, $S_{\ell}\subset\{1,\dots,n_{u}\}^{\ell}$ denotes the sets of multi-indices of the Lie brackets required for solving theES problem, $g_{I_{\ell}}$ are the corresponding Lie brackets, and $c_{I_{\ell}}$ are constant parameters. For example, for system (2) with $r=2$, $f_{0}=0$, $f_{i}=g_{i}$, $i\in\{1,\dots,n\}$, we mean $N=2$, $S_{1}=\emptyset$, $S_{2}=\{(i,j):1\leq i<j\leq n\}$, $I_{2}:=(i,j)\in S_{2}$ $c_{I_{2}}=\nu_{ij}$.
One of the main tools exploited in this paper is the Chen–Fliess series expansion [28, 17]: under certain regularity assumptions on the control vector fields of system (4), the solutions of (4) with $x(0)=x^{0}$ can be represented as
@@
$\times\int_{0}^{t}\int_{0}^{s_{1}}.\int_{0}^{s_{\ell-1}}u_{k_{1}}^{\varepsilon}(s_{1})u_{k_{\ell}}^{\varepsilon}(s_{\ell})ds_{\ell}.ds_{1}+R(t),$
@@
$R(t)=$ $\hskip-9.24994pt\sum_{k_{1},\dots,k_{N+1}=1}^{n_{u}}\int_{0}^{t}\int_{0}^{s_{1}}.\int_{0}^{s_{N}}L_{g_{k_{N+1}}}\dots L_{g_{k_{2}}}g_{k_{1}}(x(s_{N+1}))$
$\times u_{k_{1}}^{\varepsilon}(s_{1})\times\dots\times u_{k_{N+1}}^{\varepsilon}(s_{N+1})ds_{N+1}.ds_{1}.$
@@ III-A Two-input extremum seeking system
To simplify the presentation, in this section, we assume $x\in\mathbb{R}$. To steer the solutions of an extremum seeking system towards the direction of high-order Lie brackets, we refer to control approaches from nonholonomic systems theory [6, 8, 15, 29, 17, 18]. We focus here on the two-input systems of form (4):
@@
Suppose also that $u_{1}^{\varepsilon}(t)=\varepsilon^{1/N-1}v_{1}(t/\varepsilon)$, $u_{2}^{\varepsilon}(t)={\varepsilon^{1/N-1}}v_{2}(t/\varepsilon)$, with the dithers $v_{1}(t/\varepsilon),v_{2}(t/\varepsilon)$ exciting the Lie bracket $g_{I_{N}}(z)=\big[\big[\dots[g_{1},\underbrace{g_{2}],g_{2}\big],\dots,g_{2}}_{N-1\text{ times}}]\dots]\big](z)$ at time $t=\varepsilon$, with $I_{N}=(1,\underbrace{2,\dots,2}_{N-1\text{ times}})$, in the sense that the Chen–Fliess series expansion (6) takes the form
@@
and all the other Lie brackets of length from 1 to $N$ do not appear in the above expansion. One way to construct such inputs is described in [29], other approaches can be found in, e.g., [6, 8, 13]. The concrete examples are as follows, as mentioned in [18].
@@
  • the inputs $v_{1}(t/\varepsilon)=2\sqrt{\kappa_{12}\pi}\cos\left({2\kappa_{12}\pi t}/{\varepsilon}\right)$,
    $v_{2}(t/\varepsilon)=2\sqrt{{\kappa_{12}\pi}}\sin\left({2\kappa_{12}\pi t}/{\varepsilon}\right)$, $\kappa_{12}\in\mathbb{Z}$,
    excite the first order Lie bracket $g_{1},g_{2}]$;
  • the inputs $v_{1}(t/\varepsilon)=-2\left({4\kappa_{122}\pi}\right)^{\frac{2}{3}}\cos\left({4\kappa_{122}\pi t}/{\varepsilon}\right)$,
    $v_{2}(t/\varepsilon)=\left({4\kappa_{122}\pi}\right)^{\frac{2}{3}}\cos\left({2\kappa_{122}\pi t}/{\varepsilon}\right)$,
    $\kappa_{112}\in\mathbb{Z}$, excite the second order Lie bracket $g_{1},g_{2}],g_{2}]$;
  • the inputs $v_{1}(t/\varepsilon)=6\left(2\kappa_{1222}\pi\right)^{\frac{3}{4}}\sin\left({6\kappa_{1222}\pi t}/{\varepsilon}\right)$,
    $v_{2}(t/\varepsilon)=2\left(2\kappa_{1222}\pi\right)^{\frac{3}{4}}\cos\left({2\kappa_{1222}\pi t}/{\varepsilon}\right)$,
@@
$g_{1}\circ J,g_{2}\circ J](x)=-\varphi(J(x))\nabla J(x),$ (10)
@@
The argumentation is similar to the proof of the practical exponential stability in [11, Theorem 3], so we only explain here how to derive exponential decay rate to a neighborhood of $x^{*}$ based on the representation (8). Given any $\delta>0$, let $D^{\prime}$ be any compact set such that $\overline{B_{\delta}(x^{*})}\subset D^{\prime}\subset D$. Similarly to the proof of [11, Theorem 3], one can show that there exists an $\varepsilon_{0}>0$ such that solutions of system (7) with $x^{0}\in B_{\delta}(x^{*})$ and $\varepsilon\in(0,\varepsilon_{0})$ are well-defined in $D^{\prime}$ for all $t\in[0,\varepsilon]$. To estimate the remainder in (11), denote $c_{Fm}=\sup_{x\in D^{\prime}}\sum_{k_{1},\dots,k_{m+1}=1}^{2}L_{g_{k_{m}}}\dots L_{g_{k_{2}}}g_{k_{1}}\circ J(x),$ $c_{u}=\max_{k=1,2}\sup_{t\in[0,\varepsilon]}|v_{k}(t/\varepsilon)|$. Then it is easy to see that $\|R(\varepsilon)\|\leq c_{R}\varepsilon^{1+\frac{1}{m}}$ with $c_{R}=c_{u}^{m+1}c_{Fm}$. The above estimate together with the representation (11), Assumption 2, and triangular inequality, implies that ∥x(ε)-x*2≤∥x0-x*2(1-2εγm)+ε1+1mσ, with $\gamma_{m}=c_{m}(\beta_{1}-\varepsilon c_{m}\beta_{2}^{2})$, $\sigma=c_{R}^{2}\varepsilon^{1+\frac{1}{m}}+2c_{R}\delta(1+\varepsilon c_{m}\beta_{2})$. Then following the argumentation of [11, Steps 3.I-4.I], we can establish the practical exponential decay rate to an arbitrary small neighborhood of $x^{*}$. ∎
@@
As in the paper [11], it is also possible to relax regularity assumption on the control vector fields. Namely, requirement of $g_{k}\circ J$, $k=1,2$, being $m$ times continuously differentiable in $D$ can be replaced with the following: $g_{k}\circ J\in C^{m}(D\setminus\{x^{*}\});\mathbb{R}$ and $L_{g_{k_{N+1}}}\dots L_{g_{k_{2}}}g_{k_{1}}\circ J\in C(D;\mathbb{R})$ for all $N\in\{1,\dots,m\}$, $k_{1},\dots,k_{m+1}\in\{1,2\}$. This relaxation is particularly important for deriving conditions for the “classical” exponential stability in the sense of Lyapunov, meaning that the trajectories of the extremum seeking system converge to the point $x^{*}$, rather than merely to its neighborhood. Another important condition for achieving classical exponential stability is the property of vanishing amplitudes, which requires that $g_{k}\circ J\to 0$ as $x\to x^{*}$ (see [11, Theorem 3, Part II]). Since selecting such vector fields becomes increasingly challenging in the case of higher-order Lie brackets, we leave this task, along with a rigorous formulation of the corresponding exponential stability properties, for further research.
@@ III-B Alternative design approaches
with $u_{k}^{\varepsilon}(t)=\varepsilon^{-3/4}v_{k}^{\varepsilon}(t)$, $k=1,2,3$, $v_{k}(t/\varepsilon)$ exciting the Lie bracket $g_{(1,2,3)}=[[g_{1},g_{2}],g_{3}]$, and $g_{1},g_{2},g_{3}$ such that
@@
Let $\varphi_{2}\in C^{1}(\mathbb{R};\mathbb{R})$ be any given function, the functions $g_{1},g_{2}\in C^{1}(\mathbb{R};\mathbb{R})$ satisfy (10) with $\varphi=\varphi_{2}$, and let $g_{3}=-\varphi_{2}$. Then, for all $x\in D$, [[g_1∘J,g_2∘J],g_3∘J](x)=-φ_2^2((J(x)))J^(2)(x).
@@
With the exception of special cases (e.g., Newton-based ES as in [12]), in general, the even-order derivatives are rather not helpful for classic extremum seeking problems. The goal of considering this case here is to illustrate alternative approaches to designing extremum seeking systems, both in terms of selecting control vector fields and determining the number of dithers. The formula presented in Lemma 3 can be easily extended to higher-order scenarios. For example, given a four-input system (4) with $n_{u}=4$ and any function $\varphi_{3}\in C^{1}(\mathbb{R};\mathbb{R}^{+})$ such that $\sqrt{\varphi_{3}}\in C^{1}(\mathbb{R}^{+};\mathbb{R}^{+})$, let the functions $g_{1},g_{2},g_{3}$ be chosen as in Lemma 3 with $\varphi_{2}=\sqrt{\varphi_{3}}$. Then $\big[\big[[g_{1}\circ J,g_{2}\circ J],g_{3}\circ J\big],g_{4}\circ J\big](x)=-\varphi_{3}^{2}J^{(3)}(x)$.
@@
The proof mostly repeats the proof of Theorem 1, additional details will be provided in the extended version of the paper.